OpenStudy (anonymous):

A culture started with 6,000 bacteria. After 3 hours, It grew to 7,200 bacteria. Predict how many bacteria will be present after 19 hours. Round your answer to the nearest whole number.

OpenStudy (anonymous):

what is \(7200\div 6000\) ?

OpenStudy (anonymous):

N(t) = No exp(kt) k t = ln[N(t)/No] use known values at t=3 to get k from which you can then substitute t=19 and get answer.

OpenStudy (anonymous):

oh it is not nearly that complicated lets use only the numbers you see

OpenStudy (anonymous):

\[7200\div 6000=\frac{6}{5}=1.2\] so you can use \[6000\times (1.2)^{\frac{t}{3}}\] as a model

OpenStudy (anonymous):

you want the value when \(t=19\) whip out a calculator and compute \[6000\times (1.2)^{\frac{19}{3}}\]

OpenStudy (anonymous):

What is the starting equation in your approach, @satellite73 ?